# Resource Management and Circuit Scheduling for Distributed Quantum Computing Interconnect Networks

### v3

@@ Abstract
Distributed quantum computing (DQC) has emerged as a promising approach to overcome the scalability limitations of monolithic quantum processors in terms of computing capability. However, realising the full potential of DQC requires effective resource management and circuit scheduling. This involves efficiently assigning each circuit to an optimal subset of quantum processing units (QPUs), based on factors such as their computational power and connectivity. In heterogeneous DQC networks with arbitrary topologies and non-identical QPUs, this becomes a complex challenge. This paper addresses resource management in such settings, with a focus on computing resource allocation in a quantum data center. We propose circuit scheduling and resource allocation algorithms that combine heuristic methods with a Mixed-Integer Linear Programming (MILP) formulation. Our MILP model accounts for infidelities arising from inter-QPU communication. The algorithms consider key factors including network topology, QPU characteristics, and quantum circuit structure to make efficient scheduling and allocation decisions. Simulation results demonstrate that our approach significantly improves circuit execution time and resource utilisation—measured by makespan, throughput, and QPU usage—while also reducing inter-QPU communication, compared to a baseline random allocation strategy. This work provides valuable insights into resource management strategies for scalable and heterogeneous DQC systems.
@@ I Introduction
Quantum computing has gained attention as a solution for tackling intractable problems, due to its capacity to solve them significantly faster than traditional computers. In recent years, there have been notable advancements in quantum hardware and control systems, leading to the development of noisy intermediate-scale quantum processing units (QPUs). However, despite these efforts, current quantum processors still remain limited in their computational power. In this context, distributed quantum computing (DQC) can offer a substantial benefit in scaling up quantum computing beyond the constraints of monolithic systems [1, 2, 3, 4].
Distributed quantum computing aims to harness the collective power of multiple interconnected quantum processors, enabling the execution of larger and more complex quantum algorithms. In DQC, quantum algorithms are partitioned and executed across a network of quantum processors, which are interconnected through both quantum and classical communication channels. By distributing the computational workload among multiple quantum nodes, this approach facilitates the development of scalable quantum computing systems that can surpass the limitations imposed by individual quantum processors.
@@
Distributed quantum computing is expected to progress through stages of increasing scale and heterogeneity [3]. This progression spans from the integration of multiple quantum processors within a single large quantum computer to the establishment of interconnected quantum processors across various quantum data centres. This work focuses on DQC within a single quantum data centre, where multiple quantum computers are interconnected via short-to-medium range links. A quantum data centre, in this context, refers to a facility housing quantum computers and the infrastructure for their operation and maintenance.
@@
One of the fundamental requirements for DQC is the ability to perform quantum operations between distant qubits located on separate QPUs. Throughout this paper, we refer to such operations as remote gates. Most approaches proposed in the literature for realising remote gates rely on three key components: remote entanglement generation, local quantum operations on individual QPUs, and classical communication [1, 6]. As an illustrative example, Figure 1 shows the circuit diagram for implementing a controlled-NOT (CNOT) gate based on this approach between two computing qubits, denoted by $q_{\rm cp1}$ and $q_{\rm cp2}$, belonging to separate QPUs (${\rm QPU}_{1}$ and ${\rm QPU}_{2}$). The initial step in this method involves generating entanglement between two dedicated communication qubits, denoted by $q_{\rm cm1}$ and $q_{\rm cm2}$ in Fig. 1, which facilitate the quantum communication between the QPUs. As depicted in the figure, the X operation on ${\rm QPU}_{2}$ and the Z operation on ${\rm QPU}_{1}$ are conditioned on the measurement outcome transmitted classically from the other QPU. The same approach can be extended to implement any controlled-unitary gate.
The integration of quantum networking, classical networking, and quantum computation within a DQC interconnect network requires efficient orchestration of various components and tasks. A critical element in this orchestration is quantum compilation, which translates a high-level description of a quantum program into a set of instructions to be applied to the physical hardware [7, 5]. This translation is performed through several layers of subroutines forming a compilation stack, as depicted in Fig. 2 [5]. In the context of DQC, additional tasks must be executed for an intermediate representation of the quantum program, which we refer to in this work as the Quantum Circuit (QCirc) (Fig. 2). These tasks involve various aspects of resource management, encompassing both network and compute resources. Compute resources primarily refer to the computing qubits embedded in QPUs, while communication resources include communication qubits and network components such as optical switches.
In this work, we assume that network and compute resource management are handled independently—a separation that enhances scalability and modularity, both of which are essential in the complex DQC ecosystem. This assumption is consistent with the DQC architecture described in [3]. Within this framework, we assume the network can reliably establish entanglement links as needed to support remote operations between QPUs, and that network resource management can allocate the required resources, including communication qubits, on demand. As a result, our focus in this work is limited to compute resource management.
In the framework considered in this work, compute resource management is organised into two stages: circuit scheduling and circuit partitioning. Circuit scheduling manages the queue of QCircs, determines the set of QPUs allocated to each QCirc, and addresses the QCirc-to-QPU mapping problem in the presence of multiple concurrent user requests. Circuit partitioning then optimises the division of each QCirc into smaller sub-circuits, taking into account the number of partitions, the specific QPUs assigned during scheduling, and the structure of the circuit.
This paper focuses on the circuit scheduling problem in a quantum data centre environment within the DQC framework described above. Task scheduling in classical cloud data centres is a well-studied topic in the literature, typically focusing on criteria such as maximising resource utilisation, minimising makespan, and balancing the load across computing resources [8, 9, 10, 11]. However, quantum computing introduces additional critical considerations. In particular, remote gate operations introduce infidelities that can adversely affect quantum computation tasks. These infidelities may arise from qubit decoherence during remote gate execution, as well as from imperfect remote entanglement generation between QPUs. Quantum-specific constraints, such as limited qubit coherence time, imply that circuit scheduling can directly impact not only execution time but also computational fidelity. Furthermore, prior studies [12] have shown that quantum circuits exhibit varying levels of sensitivity to errors arising from distributed execution. Therefore, a scheduling policy that accounts for these quantum-specific challenges can significantly enhance both system performance and overall computational fidelity.
Our main contributions are as follows:
  • We address the problem of circuit scheduling in DQC interconnect networks, considering both traditional resource efficiency metrics and quantum-specific challenges such as fidelity degradation arising from remote operations.
  • We consider a general heterogeneous network model that includes QPUs with varying capacities, arbitrary network topologies, and diverse types of QCircs.
  • We propose two circuit scheduling algorithms that leverage both heuristic approaches and mixed-integer linear programming (MILP) formulations. The first method is based on dynamic batch scheduling, while the second adopts an online scheduling strategy.
  • The proposed methods are extensively analysed and evaluated through detailed simulations.
I-A Related work
In the literature, extensive research has focused on quantum compilation and circuit partitioning, with fewer efforts on execution management. Given the strong interdependencies among these tasks, we present key related works on these topics in what follows.
@@ I-A Related work
Several works have explored quantum compilation for DQC. Ferrari et al. [13] discussed challenges in compiler design for DQC and analytically characterised the overhead introduced by remote gates. Cuomo et al. [14] proposed compilation techniques to optimise circuit execution time and distributed entangled state usage. They modeled time as additional circuit depth layers due to entanglement generation, but their approach assumed uniform entanglement latency and neglected network topology. In a subsequent work, Ferrari et al. [15] presented a modular compilation framework incorporating network and QPU constraints. While this framework effectively considers network configuration and QPU characteristics, it primarily focuses on compilation and partitioning for a single circuit assigned to a fixed number of QPUs.
@@
The problem of quantum circuit partitioning has been addressed in multiple research works. Daei et al. [16] represent QCircs as undirected graphs where qubits are nodes and edge weights correspond to the number of two-qubit gates shared between them. They then employ the Kernighan-Lin (K-L) algorithm to partition the graph, minimising the number of edges cut across partitions. In other studies [17, 18, 19, 20], methods such as hypergraph partitioning, bipartite graph partitioning, and genetic algorithms have been employed to minimise the number of remote gates. Andres et al. [21] extended previous works on circuit partitioning to the case of heterogeneous networks with arbitrary topologies. While these studies have focused on optimally partitioning a single QCirc assuming a fixed number of partitions, they have not explicitly considered the issue of resource competition when multiple QCircs are to be concurrently assigned.
@@
Parekh et al. [22] proposed a resource allocation and scheduling algorithm based on a greedy approach that assigns QCircs and fills QPUs on a one-by-one basis. However, their algorithm does not take into account factors such as network topology and configuration, QPU decoherence properties, and the distinctive features of QCircs.
Figure 3: A high-level abstract representation of the DQC network model.
@@ II DQC Network Model
We consider a DQC network in which QPUs are interconnected via both quantum and classical links. As discussed in [23, 24], promising architectures for quantum data centers may use optical switches to dynamically establish quantum links between QPUs. Techniques such as entanglement swapping and entanglement distillation can also be employed to generate high-fidelity entanglement between nodes. In this paper, to maintain generality, we adopt an abstract model that focuses on the logical topology of the network, without specifying the underlying physical architecture or hardware implementation.
As noted in the introduction, we assume that computational and network resource management are handled independently. Accordingly, our model focuses exclusively on computing qubits within QPUs, while communication qubits and their associated management are abstracted away.
@@
A high-level representation of our DQC network model is shown in Fig. 3. In this abstract model, each quantum node is represented by $\text{QPU}_{j}$. A logical link between two QPUs, $\text{QPU}_{j_{1}}$ and $\text{QPU}_{j_{2}}$, is characterised by two key parameters for remote gate execution: the remote gate execution time, denoted by $T_{j_{1}j_{2}}$, and the fidelity of remote entanglement generation, denoted by $F_{j_{1}j_{2}}$. This model relies on several simplifying assumptions. First, we assume on-demand entanglement generation: remote entanglement generation is initiated and completed at the start of each remote gate operation, and the resulting entangled pair is used immediately. This eliminates the need for entangled-pair storage. Second, we assume fixed link parameters, meaning that both latency and fidelity values are constant for each logical link. This implicitly assumes a constant entanglement generation rate, as well as a fixed fidelity for remote entanglement generation between communication qubits.
@@
We now turn to the key assumptions and definitions related to QPUs, the quantum nodes in our DQC network model. The set of QPUs in the network is denoted by $\mathcal{QP}=\{{\text{QPU}_{j}}\}_{j=1}^{J}$, where $J$ represents the total number of QPUs in the network. The capacity of ${\rm QPU}_{j}$, defined as its total number of computing qubits, is denoted by $N_{j}$. We also introduce a dynamic parameter $n_{j}$, which represents the current availability of ${\rm{QPU}}_{j}$. Specifically, $n_{j}=N_{j}$ if ${\rm{QPU}}_{j}$ is available, and $n_{j}=0$ otherwise. We assume that all QPUs are based on the same qubit technology and exhibit nearly identical decoherence properties, characterised by a common decoherence time parameter, $T_{\text{dec}}$. It is important to note that the specific interpretation of decoherence time can vary depending on the underlying quantum technology, such as trapped ions, superconducting qubits, and others. Nonetheless, for any given QPU technology, it is possible to identify a parameter that characterises the system’s temporal limitations due to decoherence effects. To maintain platform independence, we abstract this aspect into a single representative parameter without loss of generality.
@@ III Proposed QCirc Scheduling methods
In this section, we consider the problem of QCirc scheduling for the DQC network described above. The set of input quantum circuits to be scheduled and executed is denoted by $\mathcal{QC}=\{\text{QCirc}_{m}\}_{m=1}^{M}$, where $M$ is the total number of circuits. Each circuit $\text{QCirc}_{m}$ is characterised by three parameters:
  • (a)
    the required number of qubits, denoted by $w_{m}$;
  • (b)
    the average number of two-qubit gates acting on each qubit, denoted by $g_{m}$;
  • (c)
    the standard deviation in the number of two-qubit gates per qubit, denoted by $\sigma_{m}$.
Let $g_{mq_{i}}$ denote the number of two-qubit gates involving the $i$-th qubit in $\text{QCirc}_{m}$; then $g_{m}$ is defined as the average of $g_{mq_{i}}$ over all qubits $q_{i}$, and $\sigma_{m}$ is the corresponding standard deviation [25]:
$g_{m}=\frac{1}{w_{m}}\sum_{i=1}^{w_{m}}g_{mq_{i}},\quad\sigma_{m}=\sqrt{\frac{1}{w_{m}}\sum_{i=1}^{w_{m}}(g_{mq_{i}}-g_{m})^{2}}.$
Together, the parameters $g_{m}$ and $\sigma_{m}$ capture the internal connectivity and balance of the circuit, which are critical factors in distributed quantum computing.
In the following, we propose two primary QCirc scheduling methods. The first method, termed Batch-QCirc Scheduling, is based on dynamic batch scheduling, where a group of circuits is selected and assigned to resources collectively. The second method, referred to as Single-QCirc Scheduling, assigns each circuit individually. Throughout this work, we assume a non-preemption model, meaning that once a QCirc begins execution on its allocated QPUs, it runs to completion without interruption.
@@
III-A Dynamic Batch-QCirc Scheduling

### v4

@@ Abstract
Distributed quantum computing (DQC) has emerged as a promising approach to overcome the scalability limitations of monolithic quantum processors in terms of computational capability. However, realising the full potential of DQC requires effective resource management and circuit scheduling. This involves efficiently assigning each circuit to a subset of quantum processing units (QPUs), based on factors such as their computational power and connectivity. In heterogeneous DQC networks with arbitrary connectivity topologies and non-identical QPUs, this becomes a complex challenge. This paper addresses resource management and circuit scheduling in such settings, with a focus on computing resource allocation in a quantum data center. We propose circuit scheduling algorithms based on Mixed-Integer Linear Programming (MILP). Our MILP model accounts for errors arising from inter-QPU communication. In particular, the proposed schemes consider key factors, including network topology, QPU capacities, and quantum circuit structure, to make efficient scheduling and allocation decisions. Simulation results demonstrate that our proposed algorihtms significantly improve circuit execution time and scheduling efficiency (measured by makespan and throughput), while also reducing inter-QPU communication overhead, compared to a baseline random allocation strategy. This work provides valuable insights into resource management strategies for scalable and heterogeneous DQC systems.
Index Terms:
Distributed quantum computing, Quantum networks, Resource allocation, Quantum Circuits Scheduling
@@ I Introduction
Quantum computing has gained attention as a solution for tackling intractable problems, due to its capacity to solve them significantly faster than traditional computers. In recent years, there have been notable advancements in quantum hardware and control systems, leading to the development of noisy intermediate-scale quantum processing units (QPUs). However, despite these efforts, current quantum processors still remain limited in their computational power [1, 2]. Distributed quantum computing aims to harness the collective power of multiple interconnected quantum processors, enabling the execution of larger and more complex quantum algorithms. In DQC, quantum algorithms are partitioned and executed across a network of quantum processors, which are interconnected through both quantum and classical communication channels. By distributing the computational workload among multiple quantum processors, this approach facilitates the development of scalable quantum computing systems that can surpass the limitations imposed by individual quantum processors [3, 4].
Distributed quantum computing is expected to progress through stages of increasing scale and heterogeneity [2]. This progression spans from the integration of multiple quantum processors within a single large quantum computer to the establishment of interconnected quantum processors across various quantum data centres. A quantum data centre, in this context, refers to a facility housing quantum computers and the infrastructure for their operation and maintenance at the scale of links ranging between tens and hundreds of meters, like a modern classical data centre.
One of the fundamental requirements for DQC is the ability to perform quantum operations between distant qubits located on separate QPUs, often referred to as remote gates. Most approaches for realising remote gates rely on three key components: remote entanglement generation, local quantum operations on individual QPUs, and classical communication between them [3, 5]. As an illustrative example, Fig. 1 shows the circuit diagram for implementing a controlled-NOT (CNOT) gate based on this approach between two computing qubits, denoted by $q_{\rm cp1}$ and $q_{\rm cp2}$, belonging to separate QPUs (${\rm QPU}_{1}$ and ${\rm QPU}_{2}$). The initial step in this method involves generating entanglement between two dedicated communication qubits, denoted by $q_{\rm cm1}$ and $q_{\rm cm2}$ in Fig. 1, which facilitates the quantum communication between the QPUs. As depicted in the figure, the X operation on ${\rm QPU}_{2}$ and the Z operation on ${\rm QPU}_{1}$ are conditioned on the measurement outcome transmitted classically from each other. The same approach can be extended to implement other controlled-unitary gates.
The integration of quantum networking, classical networking, and quantum computation within a DQC interconnect network requires efficient orchestration of various components and tasks. A critical element in this orchestration is quantum compilation, which translates a high-level description of a quantum program into a set of instructions to be applied to the physical hardware [6, 7]. This translation is performed through several layers of subroutines forming a compilation stack, as depicted on the left of the Fig. 2. In the context of DQC, additional tasks must be executed for an intermediate representation of the quantum program, which we refer to as the Quantum Circuit (QCirc) (see right side of Fig. 2) in this work. These tasks involve various aspects of resource management, encompassing both network and compute resources. Compute resources primarily refer to the computing qubits embedded in QPUs, while communication resources include communication qubits and network components such as optical switches.
@@
In this work, we focus on DQC within a single quantum data centre scale, where network and compute resource management are handled independently - a separation that improves scalability and modularity, both of which are essential in the complex DQC ecosystem and are consistent with the DQC architecture as described in [2]. Within this framework, the network is assumed to reliably establish entanglement links as needed to support remote operations between QPUs, and to allocate the required resources, including communication qubits, on demand. Accordingly, this work focuses on compute resource management, comprising two key stages: circuit scheduling and circuit partitioning. Circuit scheduling manages the queue of QCircs, determines the set of QPUs allocated to each QCirc, and addresses the QCirc-to-QPU mapping problem in the presence of multiple concurrent user requests. Circuit partitioning then optimises the division of each QCirc into smaller sub-circuits, taking into account the number of partitions, the specific QPUs assigned during scheduling, and the structure of the circuit.
Job scheduling in classical cloud data centres is a well-studied topic in the literature, typically focusing on criteria such as maximising resource utilisation, minimising makespan (total completion time of all tasks), and balancing the load across computing resources [8, 9, 10, 11]. However, quantum computing introduces additional critical considerations. In particular, remote gate operations introduce infidelities that can adversely affect quantum computation tasks. These infidelities may arise from qubit decoherence during remote gate execution, as well as from imperfect remote entanglement generation between QPUs. Quantum-specific constraints, such as limited qubit coherence time, imply that circuit scheduling can directly impact not only execution time but also computational fidelity. Furthermore, prior studies [12] have shown that quantum circuits exhibit varying levels of sensitivity to errors arising from distributed execution. Therefore, a scheduling policy that accounts for these quantum-specific challenges can significantly enhance both system performance and overall computational fidelity. For example, sparsely connected circuits, those with relatively few two-qubit interactions, can often be allocated with minimal inter-QPU communication. In contrast, densely connected circuits might require careful mapping to avoid excessive remote gates. If a sparse circuit is scheduled and allocated first, it may occupy the most favourable QPUs, forcing subsequent dense circuits into fragmented and suboptimal allocations. This can increase inter-QPU communication overhead, extend job execution time, and ultimately create scheduling bottlenecks at the system level. These considerations motivate the need for scheduling strategies that jointly account for multiple circuits, rather than treating them independently.
@@
While extensive research in the literature has addressed different aspects of DQC, such as network resource management, quantum compilation, and circuit partitioning, relatively fewer efforts have focused on compute resource management, particularly considering the impact of quantum circuit scheduling. Given the strong interdependencies among these tasks, we now review key related works across these areas to position our contribution in context. Resource management and scheduling in DQC have been considered in [13, 14, 15, 16]. Nevertheless, these works primarily focus on network resource management, such as scheduling entanglement requests, quantum routing, and designing efficient entanglement generation or allocation protocols. In contrast, our work addresses compute resource management, specifically job (QCirc) scheduling and QPU–QCirc mapping in heterogeneous DQC networks.
@@
Quantum compilation for DQC has been explored in several works. Ferrari et al. [17] discussed challenges in compiler design for DQC and analytically characterised the overhead introduced by remote gates. Cuomo et al. [18] proposed compilation techniques to optimise circuit execution time and distributed entangled state usage. They modeled time as additional circuit depth layers due to entanglement generation, but their approach assumes a uniform entanglement generation rate and neglects the dependency of network topology and associated loss. In a subsequent work, Ferrari et al. [19] presented a modular compilation framework incorporating network and QPU constraints. While this framework effectively considers network configuration and QPU characteristics, it primarily focuses on compilation and partitioning for a single circuit assigned to a fixed number of QPUs.
@@
Circuit partitioning has also received significant attention. Daei et al. [20] represent QCircs as undirected graphs where qubits are nodes and edge weights correspond to the number of two-qubit gates shared between them. They then employ the Kernighan-Lin (K-L) algorithm to partition the graph, minimising the number of edges cut across partitions. In other studies [21, 22, 23, 24, 25], methods such as hypergraph partitioning, bipartite graph partitioning, window-based partitioning, and genetic algorithms have been employed to minimise the number of remote gates. Andres et al. [26] extended previous works on circuit partitioning to the case of heterogeneous networks with arbitrary topologies. While these studies have focused on optimally partitioning a single QCirc assuming a fixed number of partitions, they have not explicitly considered the issue of resource competition when multiple QCircs are to be concurrently assigned.
@@
There exist gaps in DQC, particularly a lack of advanced QCirc scheduling algorithms that account for circuit partitioning when mapping and allocating resources for efficient distributed quantum computing. To the best of our knowledge, only Parekh et al. [27] proposed a greedy scheduling algorithm that assigns QCircs and fills QPUs on a one-by-one basis. However, their algorithm does not consider factors such as network topology and configuration, QPU decoherence properties, and the distinctive features of QCircs.
Here, we address the problem of QCircs scheduling in DQC networks, accounting for quantum-specific challenges such as fidelity degradation from remote operations. We adopt a general heterogeneous network model with QPUs of varying capacities, arbitrary connectivity topologies, and diverse QCircs. On this basis, we propose two scheduling algorithms: (i) a dynamic batch scheduling approach that leverages MILP to jointly optimise QPU–QCirc mapping and the number of circuit partitions for a batch of QCircs, and (ii) an MILP-based online scheduling strategy that allocates circuits individually. The proposed methods are extensively analysed and evaluated through detailed simulations.
The remainder of this paper is organised as follows. Sec. II describes the DQC network model. Section III presents the proposed Batch-QCirc and Single-QCirc scheduling methods. Section IV evaluates the proposed methods through simulations. Finally, Section V concludes the paper.
@@ II DQC Network Model
We investigate a DQC network within a quantum data centre, where QPUs are interconnected via both quantum and classical links. Promising architectures for such data centre networks will employ high-speed optical switches to dynamically establish quantum links between QPUs. Quantum information of stationary computing qubits can be encoded in photons via transducers, using degrees of freedom such as time-bin, polarization, frequency, etc [28, 29]. At the physical and link layers, remote entanglement between two QPUs can be generated in various ways [30], with one example illustrated in Fig. 3. This method relies on a Bell state measurement (BSM) module placed between the QPUs, enabling heralded entanglement generation. From an architectural perspective, switch-based data centre topologies such as fat-trees are promising candidates. A representative example with four pods is shown in Fig. 4(a), where QPUs are interconnected through three layers of optical switches to provide all-to-all connectivity. Alternatively, some designs may leverage entanglement swapping, allowing a QPU to act as a repeater and connect QPUs that are not directly linked.
Figure 3: High-level schematic of remote entanglement generation between a pair of QPUs via an intermediate BSM module. Blue circles denote computing qubits, and the brown circle denotes the communication qubit.
At a higher abstracted connectivity level, the DQC network can be represented as in Fig. 4(b), which highlights the key components: QPUs connected by quantum and classical links, coordinated by a Quantum Network Controller. This abstraction captures the essential elements of the network while remaining agnostic to specific physical implementations and architectural details.
@@
For our resource allocation and circuit scheduling problem, we further abstract the network into a logical graph model $G(V,E)$, where QPUs are represented as nodes and quantum logical links as edges. Each edge between $QPU_{j_{1}}$ and $QPU_{j_{2}}$ is characterised by two parameters: the remote gate execution time $T_{j_{1}j_{2}}$ and the fidelity of remote entanglement generation $F_{j_{1}j_{2}}$. This model relies on several simplifying assumptions. First, we assume on-demand entanglement generation: remote entanglement generation is initiated and completed at the start of each remote gate operation, and the resulting entangled pair is used immediately. This eliminates the need for long entangled-pair storage time. Second, we assume quantum communication latency and fidelity values are fixed for each logical link. In this work, we refer to the latency of a quantum communication link as the time required for one successful entanglement generation event between QPUs. Under these assumptions, remote gate execution time $T_{j_{1}j_{2}}$ is influenced by the entanglement generation rate, which typically decreases with increasing link loss. For example, in the fat-tree topology depicted in Fig. 1, link loss increases with the number of optical switches in the path, thereby increasing remote gate execution time. Link’s fidelity $F_{j_{1}j_{2}}$ is similarly influenced by factors such as link loss and the specific entanglement generation protocol used.
Figure 4: (a) Example network architecture for a DQC network within a quantum data centre (fat-tree topology with 4 pods). (b) High-level abstract representation of the DQC network model.
@@
We now turn to the key definitions and assumptions related to QPUs in our DQC network model. The set of QPUs in the network is denoted by $\mathcal{QP}=\{{\text{QPU}_{j}}\}_{j=1}^{J}$, where $J$ represents the total number of QPUs in the network. In the graph model $G(V,E)$, the vertex set $V$ corresponds to $\mathcal{QP}$, i.e., $V\equiv\mathcal{QP}$. The capacity of ${\rm QPU}_{j}$, defined as its total number of computing qubits, is denoted by $N_{j}$. We also introduce a binary parameter $s_{j}$, which represents the current availability of ${\rm{QPU}}_{j}$. Specifically, $s_{j}=1$ if ${\rm{QPU}}_{j}$ is available, and $s_{j}=0$ otherwise. We assume that all QPUs are based on the same technology and exhibit nearly identical decoherence properties, characterised by a common decoherence time parameter, $T_{\text{dec}}$. It is important to note that the concept of “decoherence time” is technology-dependent. For example, in superconducting-qubit systems, two parameters, energy relaxation time $T_{1}$ and dephasing time $T_{2}$, are often reported; these can be combined into a single effective value, such as $\min(T_{1},T_{2})$ or $\frac{T_{1}T_{2}}{T_{1}+T_{2}}$, for modelling purposes. In trapped-ion systems, a memory lifetime parameter may be used instead. Regardless of the specific definition, there always exists a parameter (or suitable combination thereof) that captures the temporal limitations imposed by decoherence. To maintain platform independence, we abstract this into a single representative parameter $T_{\text{dec}}$ in our model.
@@ III Proposed QCirc Scheduling methods
In this section, we consider the problem of QCirc scheduling for the DQC network described above. Our primary objective is to schedule a set of quantum circuits and allocate a subset of QPUs to each of them. The set of input quantum circuits is denoted by $\mathcal{QC}=\{\text{QCirc}_{m}\}_{m=1}^{M}$, where $M$ is the total number of circuits. The required number of qubits for each circuit $\mathcal{QC}=\{\text{QCirc}_{m}\}$ is represented by $w_{m}$. We define $K_{m}^{\max}$ as the maximum allowable number of partitions for $\text{QCirc}_{m}$, determined according to factors such as data centre policy, circuit size, or error tolerance. Throughout this work, we assume a non-preemption model, meaning that once a QCirc begins execution on its allocated QPUs, it runs to completion without interruption. Furthermore, each QPU can host at most one partition at any given time, meaning it cannot concurrently execute partitions from different circuits. In the following, we propose two primary QCirc scheduling methods. The first method, termed Batch-QCirc Scheduling, is based on dynamic batch scheduling, where a group of circuits is selected and assigned to QPUs collectively. The second method, referred to as Single-QCirc Scheduling, assigns each circuit individually.
@@
III-A Dynamic Batch QCirc Scheduling
To efficiently schedule diverse quantum circuits (e.g. quantum Fourier transform), with varying qubit requirements and structures, we propose a batch scheduling framework that jointly optimises QPU allocation and the number of partitions for a batch of circuits. The algorithm operates in repeated scheduling cycles, each consisting of four main stages: (a) batch selection, where incoming quantum circuits are grouped into a batch based on their combined qubit requirements; (b) batch assignment, where the selected batch is mapped to available QPUs via a holistic MILP-based optimisation that accounts for decoherence effects and inter-QPU connectivity; (c) overflow handling, where any circuits that could not be assigned in the main step are promptly scheduled to run as soon as resources become available; and (d) filling idle QPUs, where any QPUs that remain unallocated after the batch assignment process are identified and, if suitable candidates exist in the remaining circuits of $\mathcal{QC}$, assigned to them, enhancing the utilisation of available resources.